# Unsupervised Domain Adaptation with an Unobservable Source Subpopulation

### v1

@@ 1 Introduction
Unsupervised domain adaptation (UDA) (Kouw & Loog, 2019) addresses the challenge of transferring predictive models from a labeled source domain to an unlabeled target domain under distributional shifts (Koh et al., 2021; Sagawa et al., 2022). In this area, research methods aim to reduce domain discrepancy by aligning feature distributions, using statistical measures such as maximum mean discrepancy (MMD) (Tzeng et al., 2014) and higher-order moment matching (HoMM) (Chen et al., 2020). Deep adaptation frameworks, such as deep adaptation network (DAN) (Long et al., 2015) and domain-adversarial neural network (DANN) (Ganin et al., 2016), are also popularly used due to their strong empirical performance. There are also other approaches that integrate reconstruction objectives to disentangle domain-invariant and domain-specific components (Ghifary et al., 2016). These approaches often assume access to a representative and diverse set of source examples. However, real-world datasets may violate this assumption in systematic and non-random ways.
In this work, we focus on a more challenging and practically relevant UDA setting where a structured subpopulation is entirely missing from the source domain. Specifically, we consider binary label $Y$ and a binary background or environment variable $A$, and study the case where one subpopulation—defined by a particular combination of $Y$ and AA—is unobserved in the source. This structured missingness is not merely a sampling artifact, but often reflects real-world constraints in data collection. For instance, in the widely studied Waterbirds dataset (Sagawa et al., 2019), waterbirds ($Y=1$) photographed in water environments ($A=1$) can be rare or entirely absent due to the difficulty of capturing such images in the wild. This issue arises in many other disciplines as well. In healthcare, certain patient subgroups—defined jointly by disease status and demographics—may be underrepresented or absent in historical datasets due to restrictive inclusion criteria or changes in clinical practice over time. When such models are applied to broader populations, unobserved subgroups can suffer from systematic mispredictions. This structured missingness (Mitra et al., 2023) fundamentally changes some statistical properties when comparing the source and target domains, and, if unaddressed, can lead to severely biased estimation and unreliable prediction in the target domain. These structured gaps pose new challenges that are not adequately addressed by conventional UDA techniques, which motivates our work.
@@ 2 Problem Setup and Notation
Source
1
1
1
0
1
0
1
0
0
$n_{101}$
$n_{110}$
$n_{100}$
$p_{101}=\alpha_{01}\pi$
$p_{110}=\alpha_{10}\pi$
$p_{100}=\alpha_{00}\pi$
$\xi_{1}({\bf x})=\hbox{pr}(Y=1\mid{\bf X}={\bf x},A=1,R=1)$
@@
Target 0 ? 1 $n_{0\cdot 1}$ $p_{0\cdot 1}=(\beta_{11}+\beta_{01})(1-\pi)$
$\eta_{1}({\bf x})=\hbox{pr}(Y=1\mid{\bf X}={\bf x},A=1,R=0)$
$\eta_{0}({\bf x})=\hbox{pr}(Y=1\mid{\bf X}={\bf x},A=0,R=0)$
$\eta({\bf x})=\hbox{pr}(Y=1\mid{\bf X}={\bf x},R=0)$
@@
Note that this assumption is made without loss of generality, as alternative combinations–such as $(Y=0,A=1)$, $(Y=1,A=0)$, or (Y=0,A=0)(Y=0,A=0)–can be similarly assumed to have zero probability. To characterize the distributional connection between the two domains, we impose a structured conditional invariance assumption:
@@
that is, the conditional distribution of features $\bf X$} given $(Y,A)$ remains the same across domains. This can be regarded as a conditional version, or, more nuanced version, of label shift where the marginal distribution of labels (now, the combination of both label and background) varies across domains (e.g., Du Plessis & Sugiyama, 2014; Garg et al., 2020; Iyer et al., 2014; Lipton et al., 2018; Nguyen et al., 2016; Tasche, 2017; Tian et al., 2023; Zhang et al., 2013). It indicates, conditional on background $A$, the label shift assumption holds. Mathematically, it states that $p({\bf X}|Y,A,R=1)=p({\bf X}|Y,A,R=0)$. It is equivalent to $p(R|{\bf X},Y,A)=p(R|Y,A)$, the independence between $R$ and $\bf X$}, conditional on $(Y,A)$. In practice, this assumption may be suitable in many applications. Below we give two examples to illustrate the rationality of this assumption. For instance, we aim to predict user clicks on advertisements for a new batch of users (target domain, $R=0$) using historical data (source domain, $R=1$). Conditional on the advertisement type $A$ and whether the user clicks $Y$, the distribution of browsing behavior features $\bf X$} is assumed to remain stable across time periods. This is because user clicks are fundamentally determined by ad content and user interests, not by the time period in which data are collected. As another example, suppose we have datasets from two hospitals ($R=1$ indicates the source hospital and $R=0$ indicates the target hospital). Here, $\bf X$} represents imaging features, $Y$ is the disease type, and $A$ denotes patient attributes such as gender or age group. Then, conditional on the disease type $Y$ and demographic attributes $A$, the distribution of imaging features $\bf X$} is expected to remain the same across hospitals. This is because imaging characteristics for a given disease and demographic group are not systematically altered by the hospital. The main difference between hospitals lies in sampling proportions rather than in conditional distributions.
This framework captures real-world scenarios in which a certain label-background subpopulation is absent from the source domain. For example, in the Waterbirds dataset, waterbirds on water backgrounds (label $Y=1$, background $A=1$) are rarely observed–or even completely absent–in the training set, making the adaptation to target domains particularly challenging. For illustration purposes, Table 2 below shows the three observed subpopulations in the source as well as the four subpopulations in the target in two real-world datasets.
@@
Y=0:Landbird
A=1:Water background
Y=1:Waterbird
A=0:Land background
Y=0:Landbird
A=0:Land background
Y=1:Waterbird
A=1:Water background
Y=0:Landbird
A=1:Water background
Y=1:Waterbird
A=0:Land background
Y=0:Landbird
A=0:Land background
@@
Y=0:Blond hair
A=1:Male
Y=1:Dark hair
A=0:Female
Y=0: Blond hair
A=0:Female
Y=1:Dark hair
A=1:Male
Y=0:Blond hair
A=1:Male
Y=1: Dark hair
A= 0:Female
Y=0:Blond hair
A=0:Female
@@ 3.4 Estimating parameters of interest
Note that the distribution $p({\bf x}\mid R=0,A=0)$ is identifiable from the target population. The distributions $p_{10}({\bf x})$ and $p_{00}({\bf x})$ can be consistently estimated from the source population subgroups $(R=1,Y=1,A=0)$ and $(R=1,Y=0,A=0)$, respectively. Thus, the parameters ${\boldsymbol{\beta}}=(\beta_{00},\beta_{10})^{\rm T}$ can be estimated by minimizing a suitable discrepancy measure between the two sides of equation (9), such as an $L_{2}$ norm or a divergence-based criterion (e.g., Kullback–Leibler divergence), subject to the constraint in equation (10). Therefore, we reformulate the estimation of $\boldsymbol{\beta}$} as a constrained distribution matching problem:
@@
The proof of Lemma 2 is provided in Appendix S.1. A key advantage of minimizing the KL divergence is that it circumvents the need to explicitly estimate the generative models $p_{10}({\bf x})$ and $p_{00}({\bf x})$, which are often difficult to model accurately in high dimensions. Instead, it suffices to estimate one background-specific prediction model $\xi_{0}({\bf x})$ using standard classification techniques on the source domain restricted to $A=0$.
@@
6:Estimate $\xi_{1}({\bf x})$ following equation (5), as $\widehat{\xi}_{1}({\bf x})$;
@@
8:Estimate $\boldsymbol{\beta}$} and $\alpha_{y,a}$ following equations (12) and (1), as $\widehat{\boldsymbol{\beta}}$ and $\widehat{\alpha}_{ya}$ for $(y,a)\in\{0,1\}$;
9:Estimate $\eta_{1}({\bf x})$, $\eta_{0}({\bf x})$ and $\eta({\bf x})$ following equation (8), as $\widehat{\eta}_{1}({\bf x})$, $\widehat{\eta}_{0}({\bf x})$ and $\widehat{\eta}({\bf x})$.
Finally, based on all of the above discussions, we summarize the implementation details of our proposed method in Algorithm 1.
@@ 4 Theoretical Results
To establish our generalization bound, we utilize the concept of Rademacher complexity (Bartlett & Mendelson, 2002), denoted as $\mathcal{R}_{n}(\mathcal{G})$ (see Appendix S.2 for details), and impose the following assumption on the loss function:
@@ 5 Experiments
We conduct extensive experiments on both synthetic and real-world datasets. Due to space limitations, we present the results on the Waterbirds dataset here and provide additional results in Appendix S.3.
The Waterbirds dataset (Sagawa et al., 2019) consists of 11,788 images. The label $Y=1$ denotes a waterbird and $Y=0$ landbird. The background $A=1$ corresponds to a water background and $A=0$ a land background. It yields four label–background subpopulations, as summarized in Table 3.
@@
For empirical evaluation, we fix the subclass sampling rate at $a=0.5$ in the source domain and systematically vary the remaining subclass inclusion rates by setting either $b=0.5$ with $c\in\{0.1,0.2,\ldots,0.9\}$, or $c=0.5$ with $b\in\{0.1,0.2,\ldots,0.9\}$. For each configuration, the data generation process is repeated 50 times to account for sampling variability. We assess performance using two widely adopted classification metrics: accuracy and F1 score. Figure 1 presents boxplots summarizing the distribution of these metrics across repetitions for our proposed estimators $\widehat{\eta}_{1}({\bf x})$ and $\widehat{\eta}({\bf x})$, alongside benchmark estimators $\widehat{\xi}_{1}({\bf x})$ and $\widehat{\xi}({\bf x})$. The results demonstrate that the proposed methods consistently outperform the benchmarks in both accuracy and F1 score, suggesting enhanced robustness to structured subpopulation missingness in the unsupervised domain adaptation setting. Due to the space limitations, additional results are provided in Appendix S.4.
6 Related Work
Out-of-distribution (OOD) generalization
OOD detection
OOD detection is the task of identifying inputs at test time that do not come from the same distribution as the training data. Its goal is to prevent a model from making confident but incorrect predictions on unfamiliar or anomalous inputs by flagging them as OOD. There are a variety of techniques developed for OOD detection in the literature. For example, (Hendrycks & Gimpel, 2017) introduced a simple yet effective method for detecting both misclassified and OOD inputs in neural networks. Liang et al. (2018) (ODIN) proposed an improved method for detecting OOD inputs by applying temperature scaling to the softmax outputs and adding small input perturbations during inference. ODIN significantly outperformed previous baseline methods, including the maximum softmax probability approach, and set a new standard for OOD detection in classification tasks. Other techniques include but not limited to, outlier exposure (Hendrycks et al., 2018; Papadopoulos et al., 2021), ConfGAN (Sricharan & Srivastava, 2018) and OodGAN (Marek et al., 2021). In addition, Fort et al. (2021) provided an extensive empirical study of OOD detection methods across a wide range of datasets, architectures, and training regimes.
Spurious correlation
Spurious correlation is a major obstacle to OOD generalization, where models often rely on non-causal features that can degrade performance, particularly when these correlations do not generalize across domains. For example, a model trained to classify cows might rely on green pastures (background) instead of the cow itself. On a desert background, it fails. This is also the case in the Waterbirds dataset where the spurious correlation exists between label $Y$ and background $A$. Different learning strategies were proposed to discover and mitigate the impact of spurious correlation on model performance, as well as to improve model robustness. For example, (Wu et al., 2023) introduced an attention-based approach to automatically identify spurious concepts and apply adversarial training to reduce reliance on them. Another approach proposed by (Kumar et al., 2023) used causal regularization to detect and discourage spurious dependencies, allowing for scalable robustness across shifts. In addition, (Sagawa et al., 2020) investigated why overparameterization exacerbates spurious correlations, and (Kirichenko et al., 2022) found that retraining only the final layer on a small, balanced dataset can restore robustness against spurious correlations. Also, (Wang & Wang, 2024) developed a theoretical model to analyze the influence of spurious correlation strength, sample size, and feature noise on learning. Spurious correlations were also investigated in feature learning (Izmailov et al., 2022; Qiu et al., 2024), reinforcement learning (Ding et al., 2023), OOD detection (Ming et al., 2022), and text classification (Wang & Culotta, 2020). One can also resort to a comprehensive survey paper (Ye et al., 2024) on this topic.

### v2

# Unsupervised Domain Adaptation for Binary Classification with an Unobservable Source Subpopulation

@@ Abstract
Unsupervised Domain Adaptation for Binary Classification with an Unobservable Source Subpopulation
@@ 1 Introduction
Unsupervised domain adaptation (UDA) (Kouw and Loog, 2019) addresses the challenge of transferring predictive models from a labeled source domain to an unlabeled target domain under distributional shifts (Koh et al., 2021; Sagawa et al., 2022). In this area, research methods aim to reduce domain discrepancy by aligning feature distributions, using statistical measures such as maximum mean discrepancy (MMD) (Tzeng et al., 2014) and higher-order moment matching (HoMM) (Chen et al., 2020). Deep adaptation frameworks, such as deep adaptation network (DAN) (Long et al., 2015) and domain-adversarial neural network (DANN) (Ganin et al., 2016), are also popularly used due to their strong empirical performance. There are also other approaches that integrate reconstruction objectives to disentangle domain-invariant and domain-specific components (Ghifary et al., 2016). These approaches often assume access to a representative and diverse set of source examples. However, real-world datasets may violate this assumption in systematic and non-random ways.
In this work, we focus on a more challenging and practically relevant UDA setting where a structured subpopulation is entirely missing from the source domain. Specifically, we consider binary label $Y$ and a binary background or environment variable $A$, and study the case where one subpopulation, defined by a particular combination of $Y$ and $A$, is unobserved in the source. This structured missingness is not merely a sampling artifact, but often reflects real-world constraints in data collection. For instance, in the widely studied Waterbirds dataset (Sagawa et al., 2019), waterbirds ($Y=1$) photographed in water environments ($A=1$) can be rare or entirely absent due to the difficulty of capturing such images in the wild. This issue arises in many other disciplines as well. In healthcare, certain patient subgroups, defined jointly by disease status and demographics, may be underrepresented or absent in historical datasets due to restrictive inclusion criteria or changes in clinical practice over time. When such models are applied to broader populations, unobserved subgroups can suffer from systematic mispredictions. This structured missingness (Mitra et al., 2023) fundamentally changes some statistical properties when comparing the source and target domains, and, if unaddressed, can lead to severely biased estimation and unreliable prediction in the target domain. These structured gaps pose new challenges that are not adequately addressed by conventional UDA techniques, which motivates our work.
@@
2 Related Work
Out-of-distribution (OOD) generalization
OOD detection
OOD detection is the task of identifying inputs at test time that do not come from the same distribution as the training data. Its goal is to prevent a model from making confident but incorrect predictions on unfamiliar or anomalous inputs by flagging them as OOD. There are a variety of techniques developed for OOD detection in the literature. For example, (Hendrycks and Gimpel, 2017) introduced a simple yet effective method for detecting both misclassified and OOD inputs in neural networks. Liang et al. (2018) (ODIN) proposed an improved method for detecting OOD inputs by applying temperature scaling to the softmax outputs and adding small input perturbations during inference. ODIN significantly outperformed previous baseline methods, including the maximum softmax probability approach, and set a new standard for OOD detection in classification tasks. Other techniques include but not limited to, outlier exposure (Hendrycks et al., 2018; Papadopoulos et al., 2021), ConfGAN (Sricharan and Srivastava, 2018) and OodGAN (Marek et al., 2021). In addition, Fort et al. (2021) provided an extensive empirical study of OOD detection methods across a wide range of datasets, architectures, and training regimes.
Spurious correlation
Spurious correlation is a major obstacle to OOD generalization, where models often rely on non-causal features that can degrade performance, particularly when these correlations do not generalize across domains. For example, a model trained to classify cows might rely on green pastures (background) instead of the cow itself. On a desert background, it fails. This is also the case in the Waterbirds dataset where the spurious correlation exists between label $Y$ and background $A$. Different learning strategies were proposed to discover and mitigate the impact of spurious correlation on model performance, as well as to improve model robustness. For example, (Wu et al., 2023) introduced an attention-based approach to automatically identify spurious concepts and apply adversarial training to reduce reliance on them. Another approach proposed by (Kumar et al., 2023) used causal regularization to detect and discourage spurious dependencies, allowing for scalable robustness across shifts. In addition, (Sagawa et al., 2020) investigated why overparameterization exacerbates spurious correlations, and (Kirichenko et al., 2022) found that retraining only the final layer on a small, balanced dataset can restore robustness against spurious correlations. Also, (Wang and Wang, 2024) developed a theoretical model to analyze the influence of spurious correlation strength, sample size, and feature noise on learning. Spurious correlations were also investigated in feature learning (Izmailov et al., 2022; Qiu et al., 2024), reinforcement learning (Ding et al., 2023), OOD detection (Ming et al., 2022), and text classification (Wang and Culotta, 2020). One can also resort to a comprehensive survey paper (Ye et al., 2024) on this topic.
@@ 3 Problem Setup and Notation
Source 1 1 1 0 1 0 1 0 0 $n_{101}$ $n_{110}$ $n_{100}$ $p_{101}=\alpha_{01}\pi$ $p_{110}=\alpha_{10}\pi$ $p_{100}=\alpha_{00}\pi$ $\xi_{1}({\bf x})=\hbox{pr}(Y=1\mid{\bf X}={\bf x},A=1,R=1)$
@@
Target 0 ? 1 $n_{0\cdot 1}$ $p_{0\cdot 1}=(\beta_{11}+\beta_{01})(1-\pi)$ $\eta_{1}({\bf x})=\hbox{pr}(Y=1\mid{\bf X}={\bf x},A=1,R=0)$ $\eta_{0}({\bf x})=\hbox{pr}(Y=1\mid{\bf X}={\bf x},A=0,R=0)$ $\eta({\bf x})=\hbox{pr}(Y=1\mid{\bf X}={\bf x},R=0)$
@@
Note that this assumption is made without loss of generality, as alternative combinations, such as $(Y=0,A=1)$, $(Y=1,A=0)$, or $(Y=0,A=0)$, can be similarly assumed to have zero probability. To characterize the distributional connection between the two domains, we impose a structured conditional invariance assumption:
@@
that is, the conditional distribution of features $\bf X$} given $(Y,A)$ remains the same across domains. This can be regarded as a conditional version, or, more nuanced version, of label shift where the marginal distribution of labels (now, the combination of both label and background) varies across domains (e.g., Du Plessis and Sugiyama, 2014; Garg et al., 2020; Iyer et al., 2014; Lipton et al., 2018; Nguyen et al., 2016; Tasche, 2017; Tian et al., 2023; Zhang et al., 2013). It indicates, conditional on background $A$, the label shift assumption holds. It is equivalent to $p(R|{\bf X},Y,A)=p(R|Y,A)$, the independence between $R$ and $\bf X$}, conditional on $(Y,A)$. In practice, this assumption may be suitable in many applications. Below we give two examples to illustrate the rationality of this assumption. For instance, we aim to predict user clicks on advertisements for a new batch of users (target domain, $R=0$) using historical data (source domain, $R=1$). Conditional on the advertisement type $A$ and whether the user clicks $Y$, the distribution of browsing behavior features $\bf X$} is assumed to remain stable across time periods. This is because user clicks are fundamentally determined by ad content and user interests, not by the time period in which data are collected. As another example, suppose we have datasets from two hospitals ($R=1$ indicates the source hospital and $R=0$ indicates the target hospital). Here, $\bf X$} represents imaging features, $Y$ is the disease type, and $A$ denotes patient attributes such as gender or age group. Then, conditional on the disease type $Y$ and demographic attributes $A$, the distribution of imaging features $\bf X$} is expected to remain the same across hospitals. This is because imaging characteristics for a given disease and demographic group are not systematically altered by the hospital. The main difference between hospitals lies in sampling proportions rather than in conditional distributions.
This framework captures real-world scenarios in which a certain label-background subpopulation is absent from the source domain. For example, in the Waterbirds dataset, waterbirds on water backgrounds (label $Y=1$, background $A=1$) are rarely observed, or even completely absent, in the training set, making the adaptation to target domains particularly challenging. For illustration purposes, Table 2 below shows the three observed subpopulations in the source as well as the four subpopulations in the target in two real-world datasets.
@@
Y=0:Landbird A=1:Water background [Uncaptioned image] Y=1:Waterbird A=0:Land background [Uncaptioned image] Y=0:Landbird A=0:Land background [Uncaptioned image]
Y=1:Waterbird A=1:Water background [Uncaptioned image] Y=0:Landbird A=1:Water background [Uncaptioned image] Y=1:Waterbird A=0:Land background [Uncaptioned image] Y=0:Landbird A=0:Land background [Uncaptioned image]
@@
Y=0:Blond hair A=1:Male [Uncaptioned image]     Y=1:Dark hair A=0:Female [Uncaptioned image]    Y=0: Blond hair A=0:Female [Uncaptioned image]
Y=1:Dark hair A=1:Male [Uncaptioned image] Y=0:Blond hair A=1:Male [Uncaptioned image] Y=1: Dark hair A= 0:Female [Uncaptioned image] Y=0:Blond hair A=0:Female [Uncaptioned image]
@@ 4.4 Estimating parameters of interest
Note that the distribution $p({\bf x}\mid R=0,A=0)$ is identifiable from the target population. The distributions $p_{10}({\bf x})$ and $p_{00}({\bf x})$ can be consistently estimated from the source population subgroups $(R=1,Y=1,A=0)$ and $(R=1,Y=0,A=0)$, respectively. Thus, the parameters ${\boldsymbol{\beta}}=(\beta_{00},\beta_{10})^{\rm T}$ can be estimated by minimizing a suitable discrepancy measure between the two sides of (9), such as an $L_{2}$ norm or a divergence-based criterion (e.g., Kullback–Leibler divergence), subject to the constraint in (10). Therefore, we reformulate the estimation of $\boldsymbol{\beta}$} as a constrained distribution matching problem:
@@
The proof of Lemma 2 is provided in Appendix S.1. A key advantage of minimizing the KL divergence is that it circumvents the need to explicitly estimate the generative models $p_{10}({\bf x})$ and $p_{00}({\bf x})$, which are often difficult to model accurately in high dimensions. Instead, it suffices to estimate one background-specific prediction model $\xi_{0}({\bf x})$ using standard classification techniques on the source domain restricted to $A=0$. Finally, based on all of the above discussions, we summarize the implementation details of our proposed method in Algorithm 1.
@@
6:Estimate $\xi_{1}({\bf x})$ following (5), as $\widehat{\xi}_{1}({\bf x})$;
@@
8:Estimate $\boldsymbol{\beta}$} and $\alpha_{y,a}$ following (12) and (1), as $\widehat{\boldsymbol{\beta}}$ and $\widehat{\alpha}_{ya}$ for $(y,a)\in\{0,1\}$;
9:Estimate $\eta_{1}({\bf x})$, $\eta_{0}({\bf x})$ and $\eta({\bf x})$ following (8), as $\widehat{\eta}_{1}({\bf x})$, $\widehat{\eta}_{0}({\bf x})$ and $\widehat{\eta}({\bf x})$.
@@ 5 Theoretical Results
Remark 1.
Note that the tail bound described in Assumption 1 is intended to hold uniformly for every $n_{1\cdot 0}$ when estimating $\widehat{f}_{k}$ for $k=0,1$. In other words, for each subsample size $n_{1\cdot 0}$, we have a $r_{n_{1\dot{0}}}$ such that the corresponding estimators $\widehat{f}_{k}$ for $k=0,1$ are required to satisfy the stated concentration inequality. This inequality is analogous to Hoeffding’s inequality and provides a non-asymptotic concentration bound on the estimation error. Similar assumptions have also been adopted in recent work (e.g., Maity et al. (2022), Tsybakov and Audibert (2007)).
@@
To establish our generalization bound, we utilize the concept of Rademacher complexity (Bartlett and Mendelson, 2002), denoted as $\mathcal{R}_{n}(\mathcal{G})$ (see Appendix S.2 for details), and impose the following assumption on the loss function:
@@
Remark 2.
6 Synthetic Data Results
For model estimation, we utilize the known data-generating distribution to compute the conditional probability models required by both our proposed estimators and the benchmark method. Specifically, we calculate the five key conditional probabilities needed for implementation: $\xi_{0}({\bf x})$, $\xi({\bf x})$, $\tau_{0}({\bf x})$, $\tau_{1}({\bf x})$, and $\kappa({\bf x})$, which together determine the models $\eta_{0}({\bf x})$, $\eta_{1}({\bf x})$ and $\eta({\bf x})$ for our method, and $\xi_{0}({\bf x})$, $\xi_{1}({\bf x})$ and $\xi({\bf x})$ for the benchmark. Additionally, the parameters $\{\beta_{ya}:y=0,1;a=0,1\}$ and $\{\alpha_{ya}:y=0,1;a=0,1\}$ are treated as known. The classification threshold is set to $0.5$, consistent with the standard Bayesian decision rule; that is, a sample is classified as positive if the predicted probability exceeds $0.5$.
@@ 7 Experiments
In this section, we study the Waterbirds dataset (Sagawa et al., 2019), which consists of 11,788 images. This public dataset is widely used to investigate spurious correlations in image classification. It is also well aligned with our problem setting of unsupervised domain adaptation under structured missingness, where specific combinations of labels and backgrounds are systematically absent in the labeled source domain, while labels are entirely unobserved in the target domain. The label $Y=1$ denotes a waterbird and $Y=0$ landbird. The background $A=1$ corresponds to a water background and $A=0$ a land background. It yields four label–background subpopulations, as summarized in Table 3.
@@
For empirical evaluation, we fix the subclass sampling rate at $a=0.5$ in the source domain and systematically vary the remaining subclass inclusion rates by setting either $b=0.5$ with $c\in\{0.1,0.2,\ldots,0.9\}$, or $c=0.5$ with $b\in\{0.1,0.2,\ldots,0.9\}$. For each configuration, the data generation process is repeated 50 times to account for sampling variability. We assess performance using two widely adopted classification metrics: accuracy and F1 score. Figure 3 presents boxplots summarizing the distribution of these metrics across repetitions for our proposed estimators $\widehat{\eta}_{1}({\bf x})$ and $\widehat{\eta}({\bf x})$, alongside benchmark estimators $\widehat{\xi}_{1}({\bf x})$ and $\widehat{\xi}({\bf x})$. The results demonstrate that the proposed methods consistently outperform the benchmarks in both accuracy and F1 score, suggesting enhanced robustness to structured subpopulation missingness in the unsupervised domain adaptation setting. Additional results are provided in Appendix S.3.
8 Discussions and Conclusions